OpenStudy (anonymous):

Find the value of the discriminant for the equation 3x^2 - 2x + 7 = 0. -84 -76 -80 -88

OpenStudy (anonymous):

\[b^2-4ac\] \[b=-2,c=7\]

OpenStudy (anonymous):

oh and \(a=3\) compute \[(-2)^2-4\times 3\times 7\]

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

I got it wrong.. -336..

OpenStudy (anonymous):

wow!!

OpenStudy (anonymous):

\[4-84=-80\]

OpenStudy (anonymous):

It could be the calc to be honest it's older than a whale

OpenStudy (anonymous):

Thats what I did!

OpenStudy (anonymous):

yikes you are on a computer put the calculator away http://www.wolframalpha.com/input/?i=%28-2%29^2-4*3*7

OpenStudy (anonymous):

Thank you.. :(

OpenStudy (anonymous):

If the discriminant is -7, how many real number solutions does the quadratic equation have? 0 3 1 2

OpenStudy (anonymous):

ok ok , I got 1 for this am I right?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

hold on one sec

OpenStudy (anonymous):

no you are not right if the discriminant is negative there are no real solutions

OpenStudy (anonymous):

so it's 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

yw